Your company is considering adding a new production line in order to meet increasing product demand. The investment will be financed at an interest rate of 8% compounded annually. The new line will cost $650,000 in procurement and installation. An additional $21,000 will be incurred as annual operating and maintenance (O & M) cost. Additional annual revenue by using the new line is estimated to be $112,000. If the company wants to “break-even” by investing in the extra production line in 10 years. What advice will you give the company with this investment? What could be altered to make the investment more attractive?
Answer #1

We need to calculate the NPV of the investment:

  • Annual CF = Additional revenue-O&M costs
    • Annual CF = 112000-21000 = 91000
  • To break even the NPV should = 0:
Year CF Discount Factor Discounted CF
0 $ -6,50,000.00 1/(1+0.08)^0= 1 1*-650000= -6,50,000.00
1 $       91,000.00 1/(1+0.08)^1= 0.925925926 0.925925925925926*91000=        84,259.26
2 $       91,000.00 1/(1+0.08)^2= 0.85733882 0.857338820301783*91000=        78,017.83
3 $       91,000.00 1/(1+0.08)^3= 0.793832241 0.79383224102017*91000=        72,238.73
4 $       91,000.00 1/(1+0.08)^4= 0.735029853 0.735029852796453*91000=        66,887.72
5 $       91,000.00 1/(1+0.08)^5= 0.680583197 0.680583197033753*91000=        61,933.07
6 $       91,000.00 1/(1+0.08)^6= 0.630169627 0.630169626883105*91000=        57,345.44
7 $       91,000.00 1/(1+0.08)^7= 0.583490395 0.583490395262134*91000=        53,097.63
8 $       91,000.00 1/(1+0.08)^8= 0.540268885 0.540268884501976*91000=        49,164.47
9 $       91,000.00 1/(1+0.08)^9= 0.500248967 0.500248967131459*91000=        45,522.66
10 $       91,000.00 1/(1+0.08)^10= 0.463193488 0.463193488084684*91000=        42,150.61
NPV = Sum of all Discounted CF      -39,382.59

NPV is negative so this investment is not worth investing at 8%

If the interest rate is lower than 8%, the Negative NPV will improve. We need to find the rate at which NPV = 0, which is the IRR of the company. IRR can be calculated using either a financial calculator or excel or through hit and trial. Using Excel we get the IRR = 6.64% rounded to two decimal places

Year CF Discount Factor Discounted CF
0 $ -6,50,000.00 1/(1+0.0663732594891085)^0= 1 1*-650000= -6,50,000.00
1 $       91,000.00 1/(1+0.0663732594891085)^1= 0.937757948 0.937757948355806*91000=        85,335.97
2 $       91,000.00 1/(1+0.0663732594891085)^2= 0.87938997 0.879389969704491*91000=        80,024.49
3 $       91,000.00 1/(1+0.0663732594891085)^3= 0.824654934 0.824654933794758*91000=        75,043.60
4 $       91,000.00 1/(1+0.0663732594891085)^4= 0.773326719 0.773326718816865*91000=        70,372.73
5 $       91,000.00 1/(1+0.0663732594891085)^5= 0.725193277 0.725193277246431*91000=        65,992.59
6 $       91,000.00 1/(1+0.0663732594891085)^6= 0.68005576 0.680055759832036*91000=        61,885.07
7 $       91,000.00 1/(1+0.0663732594891085)^7= 0.637727694 0.637727694107639*91000=        58,033.22
8 $       91,000.00 1/(1+0.0663732594891085)^8= 0.598034214 0.598034214036059*91000=        54,421.11
9 $       91,000.00 1/(1+0.0663732594891085)^9= 0.560811338 0.560811337601031*91000=        51,033.83
10 $       91,000.00 1/(1+0.0663732594891085)^10= 0.525905289 0.525905289363419*91000=        47,857.38
NPV = Sum of all Discounted CF                  0.00